OpenStudy (anonymous):

dy/dx (ln(8x))/(9-ln(7x))

OpenStudy (anonymous):

\[ {\frac {1}{x \left( 9-\ln \left( 7\,x \right) \right) }}+{\frac { \ln \left( 8\,x \right) }{ \left( 9-\ln \left( 7\,x \right) \right) ^{2}x}} \]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Just use quotient rule.